# SG-I2V: Self-Guided Trajectory Control in Image-to-Video Generation

### v1

@@ 4.2 Qualitative Results
Figure 4: Qualitative comparison with supervised baselines. We observe that DragNUWA tends to distort objects rather than move them, and DragAnything is weak at part-level control as it is designed for entity-level control. In contrast, our method can generate videos with natural motion for diverse object and camera trajectories. Please see our project page for more comparisons.
Table 1: Quantitative comparison on the VIPSeg dataset. Despite being a zero-shot method, we show competitive motion fidelity (ObjMC) to supervised baselines without degrading video quality (FID, FVD). Furthermore, our approach outperforms other zero-shot baselines across all metrics.
@@ 4.3 Quantitative Assessment
We provide quantitative comparisons to baselines in Tab. 1. Despite being trained on large-scale datasets, Image Conductor underperforms our method by a large margin, mainly because of the limited capacity of the base model AnimateDiff. Compared to methods that also build upon SVD, we achieve competitive performance in visual quality, with slightly worse motion fidelity. Yet, these methods are trained to generate low-resolution (i.e., $320\times 576$) videos due to the high cost of fine-tuning at high resolution. At the same time, our tuning-free approach can maintain the original resolution of SVD (i.e., $576\times 1024$). We show qualitative comparisons in Fig. 4. Similar to observations in (Wu et al., 2024b), DragNUWA tends to distort objects rather than naturally move them, while our method can generate more natural movements. DragAnything is weak at part-level motion control (e.g., closing a cat’s mouth) as it is only trained on datasets annotated with entity-level control, such as object segmentation masks. In contrast, our approach can handle different granularities of control regions.
@@
The noise initialization technique in FreeTraj improves motion control slightly compared to original videos but significantly degrades visual quality. This indicates that motion prior can not be easily incorporated into initial noises of SVD by a hand-crafted algorithm. DragDiffusion leverages feature maps extracted from the upsampling block, which, unlike SG-I2V, are not semantically aligned across frames. As a result, the motion fidelity is much lower than ours. Overall, SG-I2V significantly outperforms zero-shot baselines across all metrics. This highlights that zero-shot techniques applied in text-to-video diffusion models do not necessarily transfer to image-to-video models.
@@ 4.4 Ablation Studies
Fig. 10 summarizes the effects of optimizing features from different denoising timesteps ($t=50$ corresponds to standard Gaussian noise). We first study the effect of individual timesteps. Timesteps lower than $t=30$ give significantly worse visual quality, supporting our hypothesis that only early denoising steps determine the layout of generated videos. Timesteps greater than $t=45$ also lead to bad motion fidelity, likely because detailed semantic information has yet to be generated at these extremely high noise levels. Based on these observations, we optimize feature maps extracted between timesteps 30 and 45, balancing visual quality and motion fidelity.
@@ 5 Conclusion
Appendix A Additional Qualitative Results
We strongly encourage readers to refer to our project website: https://kmcode1.github.io/Projects/SG-I2V for more visualizations in video format, where we have released qualitative results with various trajectories, comparisons with baselines, and qualitative analysis for ablation study.
Appendix B Additional Results for Ablation Study

### v2

@@ 4.2 Qualitative Results
Figure 4: Failure cases in supervised baselines. We observe that DragNUWA tends to distort objects rather than move them, and DragAnything is weak at part-level control as it is designed for entity-level control. In contrast, our method can generate videos with natural motion for diverse object and camera trajectories. Please see our project page for more general comparisons.
Table 1: Quantitative comparison on the VIPSeg dataset. Despite being a zero-shot method, we achieve small gaps in motion fidelity (ObjMC) to supervised baselines without degrading video quality (FID, FVD). Furthermore, our approach outperforms zero-shot baselines across all metrics.
@@ 4.3 Quantitative Assessment
We provide quantitative comparisons to baselines in Tab. 1. Despite being trained on large-scale datasets, Image Conductor underperforms our method by a large margin, mainly because of the limited capacity of the base model AnimateDiff. Compared to methods that also build upon SVD, we achieve competitive performance in visual quality, with slightly worse motion fidelity. Yet, these methods are trained to generate low-resolution (i.e., $320\times 576$) videos due to the high cost of fine-tuning at high resolution. At the same time, our tuning-free approach can maintain the original resolution of SVD (i.e., $576\times 1024$). Fig. 4 illustrates comparison in failure cases of supervised baselines. Similar to observations in (Wu et al., 2024c), DragNUWA tends to distort objects rather than naturally move them, while our method can generate more natural movements. DragAnything is weak at part-level motion control (e.g., closing a cat’s mouth) as it is only trained on datasets annotated with entity-level control, such as object segmentation masks. In contrast, our approach can handle different granularities of control regions.
@@
The noise initialization technique in FreeTraj improves motion control slightly compared to original videos but significantly degrades visual quality. This indicates that motion prior can not be easily incorporated into initial noises of SVD by a hand-crafted algorithm. DragDiffusion and MOFT leverages feature maps that are not semantically corresponding across frames. As a result, the motion fidelity is much lower than ours. Overall, SG-I2V significantly outperforms zero-shot baselines across all metrics. This highlights that zero-shot techniques applied in text-to-video diffusion models do not necessarily transfer to image-to-video models.
@@ 4.4 Ablation Studies
Fig. 10 summarizes the effects of optimizing features across different denoising timesteps ($t=50$ corresponds to standard Gaussian noise). Lower timesteps (e.g., $t=20,1$) significantly impair visual quality, motivating us to optimize only on earlier denoising steps. Conversely, timesteps greater than $t=45$ degrade motion fidelity due to the lack of detailed semantic information at extremely high noise levels. Notably, ObjMC improves at $t\in[20,1]$, but due to the Co-tracker tracking moving artifacts (See here). Based on these observations, we optimize feature maps extracted between timesteps 30 and 45, balancing visual quality and motion fidelity.
@@ 5 Conclusion
Appendix A Experimental Setup Details
Details of FVD. As mentioned in Sec. 4.1, we resize all the generated videos to $320\times 576$ before inputting into the evaluation scripts to ensure a fair comparison. During the computation of FVD, we further resize all the videos to $256\times 256$ following DragAnything (Wu et al., 2024c).
Details of ObjMC. Following DragAnything (Wu et al., 2024c), we compute ObjMC to evaluate motion fidelity. ObjMC is defined as the framewise average distance between the generated and ground truth trajectories, where the ground truth trajectories are reused from DragAnything and the generated trajectories are estimated using Co-Tracker v2 (Karaev et al., 2024). Since the VIPSeg dataset comprises videos with varying resolutions  (e.g., $480\times 800$, $1440\times 2560$), we resize all the videos to a uniform resolution (as specified in Tab. 1) before inputting to the diffusion model. Unlike DragAnything, which resizes the generated videos back to the original resolutions to compute ObjMC, we resize all of them to a fixed resolution of $320\times 576$. Additionally, we exclude ground truth trajectory points that are not within the image space due to objects moving out of frame. Furthermore, we exclude short videos with fewer than $14$ frames for evaluation.
Details of zero-shot adopted methods. Due to the unavailability of zero-shot trajectory-controlled image-to-video generation, we adopt techniques from text-to-image or text-to-video methods to create new baselines in Tab. 1.
FreeTraj incorporates the noise initialization technique known as trajectory injection from (Qiu et al., 2024). This work has observed that introducing motion bias into the initial noise, by copy-pasting the noise from the first frame to other frames along the trajectories, influences the motions in the generated videos of text-to-video diffusion models. However, we did not observe similar effects in image-to-video diffusion models; instead, we observed a degradation in visual quality, likely due to the latents being out of distribution.
DragDiffusion is inspired by the image editing technique in (Shi et al., 2024), which utilizes feature maps extracted from the output of upsampling blocks to guide the generation process. Concretely, we perform optimization on the feature maps output from the second and third upsampling blocks located at the middle resolution level. Although the original DragDiffusion does not include high-frequency preserved post-processing, we incorporate this in the baseline to enhance visual quality. We use the same hyperparameters as in our method.
MOFT is similar to DragDiffusion but optimizes on feature maps derived from Content Correlation Removal proposed by Xiao et al. (2024a). Concretely, we optimize on the outputs of the upsampling layers subtracted by their mean features averaged across frames. These feature maps are well-suited for reference-guided motion generation as they are motion-consistent (i.e. pixels with similar motion exhibit similar feature vectors) enabling direct comparison with reference motion feature maps. However, these feature maps are less sensitive to semantic information, and therefore, are not effective for our optimization, where reference motion is unavailable.
Appendix B Additional Qualitative Results
We strongly encourage readers to refer to our project website: https://kmcode1.github.io/Projects/SG-I2V for more visualizations in video format, where we have released qualitative results with various trajectories, qualitative results on VIPSeg dataset, qualitative comparisons with supervised and zero-shot baselines, and qualitative analysis for ablation study.
Appendix C Additional Results for Ablation Study
@@
Inference time. Our method generates $14$ frames videos at a resolution of $576\times 1024$ with $T=50$ sampling steps. Optimization is performed from the timestep $45$ to the timestep $30$ with $5$ iterations per timestep, totaling $75$ optimization iterations. The runtime depends on the number of trajectory conditions, with an average runtime of $305$ seconds on the VIPSeg dataset with A6000 48GB. Due to the need of backpropagation, the peak GPU memory usage amounts to around $30$ GB.
@@
Gaussian weighting. Following DragAnything, we weigh the feature difference using a Gaussian heatmap during the loss computation in Sec. 3.3. This accounts for potential errors in placing bounding boxes, where the bounding box may include background pixels around the object that are not intended to be moved. As shown in our ablation on the VIPSeg dataset (Tab. 2), Gaussian weighting results in small but consistent improvement across all metrics.
@@
Table 2: Ablation on Gaussian weighting on the VIPSeg dataset. Using a Gaussian heatmap in loss computation consistently improves the results across all metrics.
Mask shape FID ($\downarrow$) FVD ($\downarrow$) ObjMC ($\downarrow$)
Gaussian weighting 28.87 298.10 14.43
Identity weighting 28.88 300.20 14.72